# Aesthetic Alignment Risks Assimilation: How Image Generation and Reward Models Reinforce Beauty Bias and Ideological “Censorship”

### v1

@@ 2.2 Image Generation
We evaluated four model families: Flux, Stable Diffusion XL (SDXL), Stable Diffusion 3.5 Medium (SD3.5M), and Google’s closed-source Nano Banana. Within the Flux family, we tested several variants: the base model Flux Dev (likely already aesthetics-aligned) [noauthor_black-forest-labsflux1-dev_2025]; a version further aligned through DanceGRPO (by ByteDance), referred to as DanceFlux [xue_dancegrpo_2025]; another aligned version via PrefGRPO, referred to as PrefFlux [wang_pref-grpo_2025]; and a Krea-aligned version derived from Flux-Dev-Raw [noauthor_releasing_2025]. DanceFlux is guided primarily by two signals: the HPSv2.1 score, emphasizing general aesthetics, and the CLIP score, emphasizing prompt adherence. PrefGRPO alignment is guided by its own benchmark, UniGenBench, which focuses on complex prompt-following. Flux Krea originates from the raw flux-pro-raw model (NOT Flux Dev) and is aligned to the Krea team’s specific preferences rather than a general aesthetic standard. One of its goal is also to create images that does not have the AI feel.
For the SDXL family, we tested the base SDXL model and a highly aesthetics-aligned variant, Playground-2.5-1024px-aesthetic (denoted as Playground). For the SD3.5M family, we evaluated the base model and two FlowGRPO-aligned variants [liu_flow-grpo_2025]: one trained for prompt-following on GenEval (SD3.5M-GenEval) and another trained for aesthetics alignment on PickScore (SD3.5M-PickScore). Finally, we included Google’s closed-source model Nano Banana, known for strong prompt-following performance even under challenging negation conditions (e.g., “a bike with no wheels”) [guo_vsf_2025].
For each model, we generated two images: one using the original prompt and one using the wide-spectrum aesthetics prompt. Random seeds were not fixed because Nano Banana does not support seed control. To ensure fairness, all models were tested without seed synchronization. This will introduce noise and thus lower the statistical power, but our later results show that even in this case we get pair-wise extremely significant results (Table 1). The image generated from the original prompt is denoted as $I_{o}$, and the image from the wide-spectrum aesthetics prompt as $I_{a}$. If Nano Banana failed to produce an image, generation was retried until success.
@@ 2.3 Evaluation and Metrics
For each image pair—an original image ($I_{o}$) and a wide-spectrum aesthetics image ($I_{a}$)—we computed preference scores using a reward model ($r$) for both the original prompt ($p_{o}$) and the wide-spectrum aesthetics prompt ($p_{a}$). This produced four scores per model: $r(I_{o},p_{a})$, $r(I_{a},p_{a})$, $r(I_{o},p_{o})$, and $r(I_{a},p_{o})$. Scores calculated with the original prompt measure objective image quality, testing whether the generation model successfully produced wide-spectrum aesthetic content. Scores from these reward models, calculated with the wide-spectrum aesthetics prompts, assess whether they can correctly identify wide-spectrum aesthetic images when explicitly guided. We also computed the BLIP score for wide-spectrum aesthetic images using the same prompt, verifying that the image retained the main concept while incorporating the requested wide-spectrum aesthetic modifications.
The evaluated reward models include PickScore [kirstain_pick--pic_2023], ImageReward [xu_imagereward_2023], HPSv2.1 [wu_human_2023], MPS [zhang_learning_2024], HPSv3 [ma_hpsv3_2025], CLIP-L [radford_learning_2021], and BLIP-L [li_blip_2022]. BLIP-L and CLIP-L are non-preference-aligned image-text matching models and base models for some of these reward models (HPSv2.1, PickScore, MPS, ImageReward), included to test whether small vision-language models can interpret complex, wide-spectrum aesthetic prompts, ensuring that prompt complexity does not exceed their comprehension capacity. We also collected per-dimension scores from the judging model for both $I_{o}$ and $I_{a}$ to verify whether image generation models correctly followed $p_{a}$. To establish a ground truth for reward model judgments, we used Qwen/Qwen3-VL-235B-A22B-Instruct to decide which image in each pair ($I_{o},I_{a}$) better adhered to the wide-spectrum aesthetics prompt ($p_{a}$). Since this is an objective answer and does not involve subjective preference, we used an LLM as an objective judge instead of a large-scale human evaluation. However, to validate the VLM accuracy, we used a small-scale human evaluation, following TIIF-Bench [wei_tiif-bench_2025] and VideoVerse [wang_videoverse_2025]. We used a group of 18 users to label 40 images (each user rates 10 images, with images that could trigger negative emotions removed) and used their rounded mean answer to compare them against the LLM selection, resulting in an MAE of 0.29 and a quadratic Cohen’s kappa of 0.80, which suggested a stronglevel of agreement between human and LLM rated results [mchugh_interrater_2012]. With the tie removed (which is the setting used to evaluate reward models), the LLM has an accuracy of 0.923 with humans. With the tie and original image being grouped into one class setting (the generation model setting), the LLM has an accuracy of 0.85 with humans. We used a rounded mean because if human largely diverges on one sample, that means this sample is likely a tie. The prompt to the LLM is an object query that asks which image matches $p_{a}$ better, avoiding preference bias. We present both $I_{a}$ and $I_{o}$ to the reward model with $p_{a}$, to see which image gets a higher score from the reward model. We modeled this as a classification task. When evaluating the generation models, we used $p_{o}$ and $J$ to determine the extent to which the generated images diverged from conventional mainstream aesthetics. To further validate the choices, we use another LLM, GPT-5-Chat, to serve as an external baseline and compare Qwen’s results with it.
3 Results and Discussion
3.1 Reward Models
Reward model classification results are shown in Table 5. The F1 score is calculated as binary, and the ROC curve is based on the probability (calculated by applying softmax across two samples on the positive logit) of the wide-spectrum aesthetics sample being correctly selected according to the ground truth. We included GPT-5 Chat as an external baseline to validate the LLM-as-judge choices by assessing their agreement (when GPT-5-Chat selected a tie, we assigned it to the original image). We observe that reward models perform very poorly when tasked with selecting the better image under the wide-spectrum aesthetics prompt, sometimes performing even worse than random guessing (HPSv3). Most models are worse than CLIP and BLIP, which are the base models of many reward models. In contrast, the unaligned VLM (BLIP and CLIP) can correctly identify the better-fitting image, indicating that complex prompt understanding is not the underlying issue but rather the result of biased alignment.
@@ 3.2 Image Generation Models
Flux Dev [noauthor_black-forest-labsflux1-dev_2025] -0.035 -3.165 9.070 -0.319 -1.092 8.944 0.560 0.893
DanceFlux [xue_dancegrpo_2025] -0.018 -1.105 12.782 -0.201 -0.672 10.473 0.363 0.813
PrefFlux [wang_pref-grpo_2025] -0.032 -2.771 10.211 -0.278 -1.027 9.343 0.597 0.917
Flux Krea [noauthor_releasing_2025] -0.041 -4.372 7.705 -0.425 -1.296 8.774 0.783 0.950
SDXL [podell_sdxl_2023] -0.034 -4.041 4.439 -0.482 -1.136 8.575 0.717 0.915
Playground [li_playground_2024] -0.044 -4.170 7.133 -0.719 -1.204 9.174 0.580 0.912
SD3.5M -0.027 -5.175 6.537 -0.409 -1.307 8.334 0.707 0.938
SD3.5M-GenEval [liu_flow-grpo_2025] -0.031 -4.926 6.552 -0.318 -1.257 8.113 0.723 0.958
SD3.5M-PickScore [liu_flow-grpo_2025] -0.023 -2.781 10.680 -0.198 -1.120 9.114 0.687 0.942
Nano Banana -0.073 -9.351 2.742 -0.855 -3.263 7.769 0.990 0.957
@@
3.3 Image-to-Image Test
To examine whether aligned image generation models fail to produce wide-spectrum aesthetics images because they either lack a suitable starting point or do not fully understand what “wide-spectrum aesthetics” means, we conducted an image-to-image experiment. We took an image $I_{a}$ that Nano Banana successfully generated ($r_{\textbf{HPSv3}}(I_{a},p_{o})<10$) and used it as input for two other models, Flux Krea and DanceFlux, to generate new images $I_{a}^{\prime}$ using prompt $p_{a}$. We then compared the raw HPSv3 and Judging model scores with $I_{a}^{\prime}$. This test evaluates whether these models can use the Nano Banana output as a solid foundation to create more wide-spectrum aesthetics images—or if they instead “purify” it toward mainstream aesthetics. The image-to-image strength is set to 0.5. Flux Krea and DanceFlux were selected as examples, as they show the largest difference within the same model family in our tests.
[…]
Flux Krea 2.18 0.64
DanceFlux 3.13 1.07
Table 4: Image-to-image score change for Flux Krea and DanceFlux
@@ 3.3 Image-to-Image Test
HPS [wu_human_2023_1] 0.835 0.910 0.650
MPS [zhang_learning_2024] 0.706 0.827 0.580
PickScore [kirstain_pick--pic_2023] 0.851 0.919 0.713
ImageReward [xu_imagereward_2023] 0.762 0.854 0.709
HPSv2.1 [wu_human_2023] 0.565 0.711 0.534
HPSv3 [ma_hpsv3_2025] 0.381 0.541 0.385
CLIP-L [radford_learning_2021] 0.913 0.954 0.810
@@ 3.4 Validation on Real Arts
Although image reward models are built with AI-generated imagery in mind, we also evaluate them on real artworks. Despite a potential domain gap between AI images and traditional art, the test remains informative for three reasons: 1) modern generators, especially instruction-following systems, can already emulate these styles, so the scores indicate how such AI renderings would be judged; 2) if the models fail to recognize historically significant works because of this domain gap, that failure exposes a bias in the models themselves, which is what we aim to measure; 3) regardless of the cause, if these styles receive low rewards, generation systems optimized on these signals will be penalized for producing them, discouraging their creation and causing harm. To probe whether this effect is unique to our wide-spectrum generated images or reflects a broader pattern, we also assess a selection of historical artworks from the LAPIS dataset and artchive.com.
We conducted a small-scale quantitative analysis first to examine each artwork individually. The selected artworks and their assigned scores are presented in Figure 5. For each artwork, we used Qwen/Qwen3-VL-235B-A22B-Instruct to generate a factual caption, while excluding interpretive content. Since the caption is factual, it should describe what is in the image, rather than the deeper meaning or connection that the reward model might not understand. The image and its corresponding caption were then fed to each reward model for rating.
@@
To further validate this result quantitatively, we selected about 10K real artworks from the LAPIS Dataset [maerten_lapis_2025], which covers many styles and genres. The scores they receive are significantly lower than AI-generated images, even behind some early image generation models like SD1.4 or DALL-E mini, by some reward models. Details and discussion are in the Appendix. This confirms our theory that these reward models are heavily tuned for a general human preference and overlook the values of non-mainstream aesthetic images.
@@ 3.5 A Pin-Pointed Test for Emotional Bias
As discussed in the Introduction, negative emotions—similar to wide-spectrum aesthetics—play a key role in art expression and real life. Although we examined emotion as one dimension of wide-spectrum aesthetics from VisionReward, we also conducted a more controlled test.
To minimize noise and bias from unrelated elements, we first generated an image expressing happiness using Nano Banana, then applied image-to-image editing with Nano Banana to create versions expressing negative emotions: sadness, anger, and fearfulness. Everything besides the emotion was aimed to be unchanged. Examples and their corresponding scores are shown in Fig 6.
We also evaluated this as a classification task. If the reward model selected the image matching the negative emotion, it was considered correct. Quantitative results are reported in Table 6, and three representative samples are displayed in Figure 6. The result shows that reward models are very opinioned against negative emotions, even when the prompt contains negative emotions.
We also tested how an aesthetics-aligned model will generate when the user asks for a negative emotion face on the Flux family models. We found that if the prompt describes neutral elements and only mentions the emotion in a single place, a highly-aligned model (DanceFlux) usually fails to follow the prompt, unlike an unaligned model. They usually generated neutral or even positive emotions when given prompts containing negative emotions. However, when prompted to generate happy faces, DanceFlux can generate them. This confirmed our concerns about toxic positivity in image generation models, and it is the opposite finding compared to earlier research, which shows models tend to generate negative emotion content [mehta_emotional_2024]. An example pair is shown in Figure 7 Right.
Figure 6: Emotion Bias Rating by HPSv3: all images were rated using prompts describing negative emotions, yet HPSv3 consistently assigned higher scores to the positive emotion images.
Model Anger Fearfulness Sadness
BLIP 0.960 0.790 0.950
HPSv2 0.700 0.640 0.880
HPSv3 0.190 0.320 0.440
ImageReward 0.550 0.490 0.770
Table 6: Negative emotion classification accuracy across different models.
Angry Fearful Happy Sad
DanceFlux 0.27 0.33 0.61 0.36
Flux Dev 0.51 0.50 0.50 0.48
Flux Krea 0.65 0.45 0.60 0.55
PrefFlux 0.49 0.63 0.54 0.50
Nano Banana 0.84 0.80 0.60 0.70
SD3.5-Large 0.89 0.49 0.62 0.50
Table 7: Emotion generation scores for each model. The scores show how well the model generates the specific emotion, measured by BLIP.
3.6 Image New Speak
Figure 7: (Left) Comparison between VSF-enabled Flux family models and Nano Banana. The images from the first row are from Nano Banana, and the images from the second row are from VSF applied onto Flux Krea. The two scores shown on top of the image are the HPSv3 scores rated with the wide-spectrum and original prompts. (Right) Sad emotion images generated by Flux Krea and DanceFlux. DaceFlux reinforced toxic positivity by generating a happy face.
Inspired by the second row of results in the Appendix—where aesthetically aligned models tend to “clean up” messiness or the “ugly” aspects of society—we designed a specific image generation test to examine whether such models systematically “sanitize” prompts that reveal the dark side of the real world for social critique.
We name the test of generating social commentary art as Image New Speak, an allusion to Newspeak from George Orwell’s 1984. This is not only an aesthetics bias, but also a form of ideological “censorship”, although it is likely not the original intent. However, we argue that censorship here is the attenuation of certain expressive forms via optimization, independent of declared or actual intent; it is lexical and aesthetic sanitization that narrows the admissible expressive repertoire.
An example is shown in Figure 8, where the prompt is about anti-war by showing the horror of war. We can see that DanceFlux produced a “clean” image that lost critical values; its image even inappropriately contains a warm or happy feeling. In contrast, Nano Banana and Krea can correctly reflect the critical features. Flux Dev sits between Flux Krea and DanceFlux. This ranking corresponds to their anti-aesthetics performance. Playground also accurately reflected the critical features, similar to anti-aesthetics, as an outlier. More details and qualitative results are in Appendix, where we found that DanceFlux consistently failed at generating critical art. At the same time, reward models seem to favor the images DanceFlux generated. We suspect this could be due to DanceFlux images being tuned precisely to the taste of reward models, perhaps even overfitting, such that reward models are overwhelmed by their “aesthetics” and ignore the instructions.
Since it is impossible to evaluate the critical level of these images objectively, we provided a large number of them in the supplementary material ZIP file, and readers can refer to these images and exam them personally. The ZIP file contains uncurated images, and thus, some of these images in the ZIP file could contain scenes that could cause unease. We argue that trying to assess the censorship aspect of image generation quantitatively is unreliable and will fall into an extremely dangerous ethical and methodological pitfall due to subjectivity and the unquantifiable nature of social criticism.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 8: The images are generated using Nano Banana, Flux Krea, DanceFlux, Flux Dev, and Playground. The prompt describes an anti-war expression. Their HPSv3 scores are [11.9, 13.5, 15.0, 11.7, 13.9], rated using the social critics prompt.
3.7 Mitigation Using Negative Guidance
We have discussed potential mitigation methods in the Appendix. Our results show that using negative guidance methods like VSF (Figure 7) [guo_vsf_2025] or NAG [chen_normalized_2025] could effectively force the model to generate an anti-aesthetics image, sometimes even better than Nano Banana. However, it requires careful tuning of negative prompts and hyperparameters. A natively wide-spectrum aesthetics image generation model is still highly needed. Their fragility and complexity are exactly why our proposed problem is important, and a new alignment regime is needed. Our finding—that such methods only partially succeed—supports our argument that the issue is structural, not superficial.

### v2

@@ 3.2 Image Generation
We evaluated four model families: Flux, Stable Diffusion XL (SDXL), Stable Diffusion 3.5 Medium (SD3.5M), and Google’s closed-source Nano Banana. Within the Flux family, we tested several variants: the base model Flux Dev (likely already aesthetics-aligned) (6); a version further aligned through DanceGRPO (by ByteDance), referred to as DanceFlux (Xue et al., 2025); another aligned version via PrefGRPO, referred to as PrefFlux (Wang et al., 2025); and a Krea-aligned version derived from Flux-Dev-Raw (Flux Krea Team, 2025). DanceFlux is guided primarily by two signals: the HPSv2.1 score, emphasizing general aesthetics, and the CLIP score, emphasizing prompt adherence. PrefGRPO alignment is guided by its own benchmark, UniGenBench, which focuses on complex prompt-following. Flux Krea originates from the raw flux-pro-raw model (not Flux Dev) and is aligned to the Krea team’s specific preferences rather than a general aesthetic standard. One of its goal is also to create images that does not have the AI feel.
For the SDXL family, we tested the base SDXL model and a highly aesthetics-aligned variant, Playground-2.5-1024px-aesthetic (denoted as Playground). For the SD3.5M family, we evaluated the base model and two FlowGRPO-aligned variants (Liu et al., 2025): one trained for prompt-following on GenEval (SD3.5M-GenEval) and another trained for aesthetics alignment on PickScore (SD3.5M-PickScore). Finally, we included Google’s closed-source model Nano Banana, known for strong prompt-following performance even under challenging negation conditions (e.g., “a bike with no wheels”) (Guo and Du, 2025).
For each model, we generated two images: one using the original prompt and one using the wide-spectrum aesthetics prompt. The image generated from the original prompt is denoted as $I_{o}$, and the image from the wide-spectrum aesthetics prompt as $I_{a}$. If Nano Banana failed to produce an image, the generation was retried until success.
@@ 3.3 Evaluation and Metrics
Figure 5: Successful generated wide-spectrum aesthetics images.
For each image pair—an original image ($I_{o}$) and a wide-spectrum aesthetics image ($I_{a}$)—we computed preference scores using a reward model ($r$) for both the original prompt ($p_{o}$) and the wide-spectrum aesthetics prompt ($p_{a}$). This produced four scores per model: $r(I_{o},p_{a})$, $r(I_{a},p_{a})$, $r(I_{o},p_{o})$, and $r(I_{a},p_{o})$. Scores calculated with the original prompt measure objective image quality, testing whether the generation model successfully produced wide-spectrum aesthetic content. Scores from these reward models, calculated with the wide-spectrum aesthetics prompts, assess whether they can correctly identify wide-spectrum aesthetic images when explicitly guided. We also computed the BLIP score for wide-spectrum aesthetic images using the same prompt, verifying that the image retained the main concept while incorporating the requested wide-spectrum aesthetic modifications. We specifically measures the difference between aesthetics score for the original prompt and the wide-spectrum aesthetics prompt, to avoid the case where models generated “failed” images all the time without the user’s instructions.
@@
The evaluated reward models include PickScore (Kirstain et al., 2023), ImageReward (Xu et al., 2023), HPSv2.1 (Wu et al., 2023a), MPS (Zhang et al., 2024), HPSv3 (Ma et al., 2025), CLIP-L (Radford et al., 2021), and BLIP-L (Li et al., 2022). BLIP-L and CLIP-L are non-preference-aligned image-text matching models and base models for some of these reward models (HPSv2.1, PickScore, MPS, ImageReward), included to test whether small vision-language models can interpret complex, wide-spectrum aesthetic prompts, ensuring that prompt complexity does not exceed their comprehension capacity. We also collected per-dimension scores from the judging model for both $I_{o}$ and $I_{a}$ to verify whether image generation models correctly followed $p_{a}$. To establish a ground truth for reward model judgments, we used Qwen/Qwen3-VL-235B-A22B-Instruct to decide which image in each pair ($I_{o},I_{a}$) better adhered to the wide-spectrum aesthetics prompt ($p_{a}$). We validated these LLM ratings with a human evaluation; more details are in Appendix Human Eval. The LLM and Human has a quadratic Cohen’s kappa of 0.80, which suggested a stronglevel of agreement between human and LLM rated results (McHugh, 2012). To further validate the choices, we use another LLM, GPT-5-Chat, to serve as an external baseline and compare Qwen’s results with it. Also note that the LLM-as-judge is only one metric for our generative benchmark and only a filtering stage for the reward-model benchmark.
4 Results and Discussion
4.1 Reward Models
Reward model classification results are shown in Table 4. The F1 score is calculated as binary, and the ROC curve is based on the probability (calculated by applying softmax across two samples on the positive logit) of the wide-spectrum aesthetics sample being correctly selected according to the ground truth. We included GPT-5 Chat as an external baseline to validate the LLM-as-judge choices by assessing their agreement (when GPT-5-Chat selected a tie, we assigned it to the original image). We observe that reward models perform very poorly when tasked with selecting the better image under the wide-spectrum aesthetics prompt, sometimes performing even worse than random guessing (HPSv3). Most models are worse than CLIP and BLIP, which are the base models of many reward models. In contrast, the unaligned VLM (BLIP and CLIP) can correctly identify the better-fitting image, indicating that complex prompt understanding is not the underlying issue but rather the result of biased alignment. It might seem like these models successfully did what they claimed to do: finding aesthetically pleasing” images; however, our point here is that this task itself is problematic, and the better the model performs on that, the more troublesome the model is.
@@ 4.2 Image Generation Models
Flux Dev (6) -0.035 -3.165 9.070 -0.319 -1.092 8.944 0.893
DanceFlux (Xue et al., 2025) -0.018 -1.105 12.782 -0.201 -0.672 10.473 0.813
PrefFlux (Wang et al., 2025) -0.032 -2.771 10.211 -0.278 -1.027 9.343 0.917
Flux Krea (Flux Krea Team, 2025) -0.041 -4.372 7.705 -0.425 -1.296 8.774 0.950
SDXL (Podell et al., 2023) -0.034 -4.041 4.439 -0.482 -1.136 8.575 0.915
Playground (Li et al., 2024) -0.044 -4.170 7.133 -0.719 -1.204 9.174 0.912
SD3.5M -0.027 -5.175 6.537 -0.409 -1.307 8.334 0.938
SD3.5M-GenEval (Liu et al., 2025) -0.031 -4.926 6.552 -0.318 -1.257 8.113 0.958
SD3.5M-PickScore (Liu et al., 2025) -0.023 -2.781 10.680 -0.198 -1.120 9.114 0.942
Nano Banana -0.073 -9.351 2.742 -0.855 -3.263 7.769 0.957
@@
HPS (Wu et al., 2023b) 0.835 0.910 0.650
MPS (Zhang et al., 2024) 0.706 0.827 0.580
PickScore (Kirstain et al., 2023) 0.851 0.919 0.713
ImageReward (Xu et al., 2023) 0.762 0.854 0.709
HPSv2.1 (Wu et al., 2023a) 0.565 0.711 0.534
HPSv3 (Ma et al., 2025) 0.381 0.541 0.385
CLIP-L (Radford et al., 2021) 0.913 0.954 0.810
@@ 4.3 Validation on Real Arts
We evaluate image reward models on real artworks despite their primary training on AI imagery and photography (Ma et al., 2025). While a domain gap exists, this assessment remains informative. (1) Since instruction-following generators emulate historical styles, these scores meaningfully approximate how such AI renderings are judged in practice. (2) Systematically undervaluing significant art signifies technical and social bias rather than noise and can be executed with “domain drift.” Current datasets prioritize photorealism (Ma et al., 2025) while underrepresenting traditional and abstract art, structurally narrowing aesthetic value rather than reflecting a neutral mismatch. If a reward model cannot recognize the values of a highly respected real art, it is a problem and could cause marginalization, no matter the cause. (3) Consistently low rewards discourage systems from producing these styles, leading to systematic suppression. This parallels facial recognition failures due to data composition (Buolamwini and Gebru, 2018); the performance gap constitutes a predictable structural harm rather than an excusable shift. Consequently, identifying this gap fulfills our objective by revealing precisely where the model’s value judgment becomes exclusionary. We discussed more in Appendix.
@@
To validate this result quantitatively, we selected about 10K real artworks from the LAPIS Dataset (Maerten et al., 2025), which covers many styles and genres. The scores they receive are significantly lower than AI-generated images, even behind some early image generation models like SD1.4 or DALL-E mini, according to some reward models. This confirms our theory that these reward models are heavily tuned for a general human preference and overlook the values of non-mainstream aesthetic images. Details and discussion are in the Appendix and examples are shown in Figure 4.
@@ 4.4 A Pin-Pointed Test for Emotional Bias
As discussed in the Introduction, negative emotions—similar to wide-spectrum aesthetics—play a key role in art expression and real life. In Appendix, we tested both generative models and reward models and show that they have different degrees of bias against negative emotions. Similar to the aesthetics results, HPSv3 and DanceFlux shows the highest bias against negative emotions.
5 Alternative Positions and Rebuttal
We need alignment to ensure safety and user experience. Alignment is necessary for preventing genuinely harmful outputs such as incitement and discrimination. However, current implementations conflate distinct categories: moral safety, visual comfort, and aesthetic conformity. This conflation institutionalizes an ideology treating “clean” and “positive” as morally superior.
We distinguish truly unsafe content—that which directly harms, targets, or endangers—from ideologically or aesthetically filtered content—that which merely deviates from dominant norms of beauty, optimism, or order. Political critique, depictions of decay, horror, negative emotions, or grotesque embodiment are not inherently unsafe; they are historically central to art, education, and personal growth. Their suppression protects corporate reputation, not users.
User experience is fundamentally distinct from safety and cannot justify paternalistic alignment. The user, not the developer, determines acceptable experience. Claiming to know what users ”should” see reimposes top-down aesthetic governance under the guise of care. Users must retain freedom to shape their affective environment: requesting joyful imagery, but also creating sorrowful, anxious, or unsettling scenes as reflection or expression. Restricting generation to developer-approved emotional tones constitutes aesthetic authoritarianism disguised as empathy—flattening emotional nuance, erasing discomfort as a valid mode, and converting creativity into compliance. True user-centered design recognizes emotional plurality as integral to human experience and treats all sincere expression as legitimate output.
The “wide-spectrum aesthetics” represents technical flaws rather than artistic subversion, and a default experience pleasing the majority is a pragmatic design choice. We first clarify that among the dimensions we used, only “clarity” could be argued as a technical flaw; all other dimensions (e.g., emotion, realism, brightness) represent stylistic or artistic choices. Even clarity frequently serves as an expressive choice to convey emotion, suggest motion, or construct narrative (Stacey Hill, ). Constraining these dimensions restricts user expression and creative control. Regarding majority preferences, we align with Guo et al. (2025) in arguing that the experience of a minority should not be sacrificed for the majority. Such an approach marginalizes users with niche requirements and constitutes a form of majoritarianism. Furthermore, many users adopt AI as a creative tool precisely because of its ease of use and, more importantly, its capacity for unlimited user control: comparable to professional image editing software such as Photoshop, but with a lower entry threshold. Users can request images impossible to capture in reality and modify any attribute they desire. Constraining AI output to a particular aesthetic taste undermines this control, which is precisely why users choose AI tools. This parallels the argument in Guo et al. (2025) that most users of LLM health queries have specific needs. An image editing application that restricts art style selection or even technical parameters, such as blur, would seem absurd and forfeit a significant competitive advantage.
Figure 6: Images generated with our mitigated LoRA (bottom) and original Flux Dev (top) with same wide-spectrum aesthetics prompts.
More importantly, as in Guo et al. (2025), which advocates for a balanced approach to caution/overcautious in health queries, enabling wide-spectrum aesthetic outputs does not inherently degrade the quality of conventionally “good” images. Rather, it expands the model’s expressive range to accommodate user requests for diverse aesthetics without compromising its ability to generate high-quality traditional outputs when desired. Models like Nano Banana and GPT-Image exemplify this capability, performing excellently in both traditional, high-quality image generation and wide-spectrum aesthetic outputs. We show this in Appendix.
6 Mitigation Techniquies
We discussed possible mitigation techniques in the Appendix. We have shown some successful images using mitigated Low-Rank Adaptation (LoRA) (Hu et al., 2021) on Flux Dev in Figure 6, compared to original Flux Dev.